# Watermarking Algorithm for Large Language Models

### v1

@@ 4.4 Results & Discussion
We compared the performance of SimMark against state-of-the-art watermarking algorithms through extensive experiments. Our primary baseline was SemStamp, a sentence-level semantic watermarking method. We also included $k$-SemStamp, an improvement over SemStamp tailored to specific text domains. Results for SemStamp, $k$-SemStamp, and other methods, including KGW and SIR, were extracted directly from Hou et al. (2024).
Table 1 presents detection performance before and after paraphrasing attacks, while Table 2 provides text quality evaluation results. Our algorithm (the last two rows in each section) does not affect the text quality while being effective and consistently outperforming or matching other state-of-the-art methods, achieving the highest average performance across all paraphrasers and datasets. Notably, domain-independent SimMark surpasses our primary baseline, SemStamp, and matches or exceeds the domain-dependent $k$-SemStamp. Refer to Appendix B for qualitative examples of SimMark.
A key aspect to consider is that the fine-tuning of SemStamp and $k$-SemStamp’s embedding model on text paraphrased by Pegasus likely contributes to its improved robustness against this specific paraphraser but may introduce bias. Additionally, the results for the Reddit-TIFU dataset were only available for SemStamp and not $k$-SemStamp, likely due to the dataset’s informal diverse text style and $k$-SemStamp’s limitation for text to belong to a specific domain such as news articles or scientific writings (Hou et al., 2024).
Figure 3 presents the detection performance (in terms of ROC-AUC) of SimMark, $k$-SemStamp, and SemStamp against Paraphrase+Delete Attack under varying sentence drop probabilities, evaluated on the RealNews and BookSum datasets. The results demonstrate that SimMark consistently achieves higher performance compared to other methods, even as the sentence drop probability increases. These findings underscore the robustness of SimMark in handling more challenging real-world adversarial scenarios.
@@
Despite Hou et al. (2023, 2024) releasing their code and data, we were unable to reproduce their reported results fully. Consequently, there are minor discrepancies between our reproduction results (shown in Figure 3 for the case with zero probability) and those presented in Table 1, which were extracted directly from their paper. Notably, the results for $k$-SemStamp on the BookSum dataset differed substantially, leading us to exclude them from Figure 3. Additionally, the results reported in Table 2 (our reproduction) also show slight differences from their paper, likely due to hyperparameter details that were not explicitly documented.
@@
Regarding sampling efficiency, for BookSum dataset for instance, cosine-SimMark required an average of 7.1 samples per sentence from the LLM, compared to $k$-SemStamp and SemStamp, which averaged 13.3 and 20.9 samples, respectively (Hou et al., 2024). This demonstrates that our method not only outperforms these baselines but is also significantly (about 2-3 times) more efficient. Further analysis of this, including theoretical estimates is provided in Appendix G.
@@ 5 Conclusion
In this paper, we introduced SimMark, a similarity-based, sentence-level posthoc watermarking algorithm. Unlike existing approaches, SimMark operates without requiring access to the internal logits or architecture of the model, ensuring compatibility with a wide range of LLMs, including API-only models. By utilizing a general-purpose sentence embedding model and integrating a soft counting mechanism, SimMark combines robustness against paraphrasing with applicability to diverse domains. Furthermore, its resistance to reverse engineering ensures resilience against spoofing attacks. Experimental results show that SimMark outperforms state-of-the-art watermarking algorithms, such as SemStamp, in both efficiency and robustness to adversarial paraphrasing, representing a significant step forward in practical watermarking techniques for LLM-generated content. Due to space constraints, the discussion of the future directions is deferred to Appendix A.
Acknowledgements
@@
Impact Statement [^7] Generative AI tools, such as ChatGPT (OpenAI, 2022), were used to refine this manuscript. The authors retain full responsibility for all content presented in the paper, ensuring adherence to academic integrity and ethical research practices.
By enabling robust detection of LLM-generated text, particularly under paraphrasing attacks, SimMark addresses ethical concerns surrounding the transparency and accountability of AI-generated content. However, like any watermarking method, there are potential risks, such as falsely implicating human authors or adversaries developing more advanced techniques for spoofing attacks or bypassing detection.
We acknowledge these limitations and advocate for the responsible deployment of such tools in combination with other verification mechanisms to mitigate these risks and ensure ethical, fair deployment. The primary goal of this work is to advance research in watermarking techniques to support the responsible use of LLMs. We believe that the societal impacts and ethical considerations of our work align with those outlined in Weidinger et al. (2021).
Appendix A Future Work
Reduce Rejection Sampling Overhead. The rejection sampling process requires generating multiple candidate sentences until a valid sentence is accepted. Although our method is 2-3 times more efficient than prior approaches such as SemStamp and $k$-SemStamp, there is still a notable decrease in generation speed due to rejection sampling. Techniques like batch sampling or parallel sampling could potentially mitigate this issue, though at the expense of higher computational resource usage. Future research should focus on optimizing these methods to balance efficiency and resource requirements.
Improve Resistance to More Advanced Attacks. Although SimMark is robust against paraphrasing attacks to some extent, more advanced text transformations (e.g., sentence splitting and merging, semantic reordering) may still pose challenges. Additionally, our detection algorithm may struggle when watermarked text is mixed into a larger body of human-written content. Further research is needed to enhance resilience against a broader range of adversarial transformations.
Setting the Interval Dynamically. In our experiments, we used consistent predefined intervals across all datasets and observed consistently strong performance. Notably, we did not observe any degradation in text quality due to this interval constraint during rejection sampling (as shown in Table 2), likely because the constraint applies only to consecutive sentences. Nonetheless, slight variations in the embeddings similarity distribution of LLM-generated text across different models and domains may impact watermarking effectiveness. Adaptive strategies for setting these intervals dynamically (or pseudo-randomly) could not only improve performance but also make reverse-engineering the algorithm even more difficult.
Appendix B Examples of Watermarked Text

### v2

@@ 4.4 Results & Discussion
Figure 3: Detection performance of different watermarking methods under various paraphrasing attacks, measured by TP@FP=1% ↑ and averaged across all three datasets (RealNews, BookSum, Reddit-TIFU). Each axis corresponds to a specific paraphrasing attack method (e.g., Pegasus-Bigram), and higher values are better. Our methods, cosine-SimMark and Euclidean-SimMark, consistently outperform or match baselines across most paraphrasers, especially under more challenging conditions such as bigram-level paraphrasing.
@@
We compared the performance of SimMark against SOTA watermarking algorithms through extensive experiments. Our primary baseline was SemStamp, a sentence-level semantic watermarking method. We also included $k$-SemStamp, an improvement over SemStamp tailored to specific domains. Results for SemStamp, $k$-SemStamp, KGW, and SIR, were extracted directly from Hou et al. (2024a, b) [^15] Despite Hou et al. (2024a, b) releasing their code and data, we were unable to reproduce their reported results fully. Consequently, there are minor discrepancies between our reproduction results (shown in Figure 4 for cases with $p=0$) and those presented in Table 1 (extracted directly from their paper). Additionally, the results reported in Table 2 (our reproduction) also show slight differences from their paper, likely due to hyperparameter details that were not explicitly documented..
Table 1 presents detection performance pre and post paraphrase attacks, while Table 2 provides text quality evaluation results. To better visualize robustness across paraphrasing attacks, we aggregate TP@FP=1% scores from Table 1 across datasets and present them in a radar plot (Figure 3). Additional radar plots and a summary table comparing performance across all paraphrasing settings are provided in Appendix C (see Table 3 and Figure 5). We also conducted a small-scale A/B test to assess watermark imperceptibility. To see the results, please refer to Appendix M. Overall, our algorithm was imperceptible to the evaluators, and impacted the text quality minimally while being effective and consistently outperforming or matching other SOTA methods, achieving the highest average paraphrased performance across all datasets.
Notably, our method, SimMark (domain-independent), surpasses the primary baseline, SemStamp (domain-independent), and is on par with or exceeds $k$-SemStamp (domain-dependent). A key aspect to consider is that the fine-tuning of SemStamp and $k$-SemStamp’s embedding model on text paraphrased by Pegasus likely contributes to their higher robustness against Pegasus but may introduce bias and reduce general applicability (as shown in Table 2 of Hou et al. (2024b), $k$-SemStamp loses performance under domain shift). Additionally, the results for the Reddit-TIFU dataset were only available for SemStamp and not $k$-SemStamp, likely due to the dataset’s informal, diverse text style and $k$-SemStamp’s limitation for text to belong to a specific domain, such as news articles or scientific writings Hou et al. (2024b).
While SimMark demonstrates strong performance, we acknowledge that it is not always the best across every setting. SimMark is designed with generality, robustness, and black-box compatibility in mind. It forgoes domain-specific fine-tuning in favor of broader applicability, which may explain why it does not always outperform more narrowly optimized methods. Several factors may explain its relative underperformance in some cases:
First, some variability arises from randomness in datasets and generation. LLMs rely on pseudo-random decoding, which may introduce subtle fluctuations in outputs. More importantly, as SimMark operates at the sentence level, it cannot fine-tune token-level patterns similar to token-level methods (e.g., UW), which may subtly manipulate every token to embed statistical signals. This granularity difference can be particularly advantageous when generation length is short (e.g., a 200-token cap). In such cases, token-level methods benefit from having more embedding capacities, whereas sentence-level methods like SimMark may yield weaker statistical signals, resulting in greater performance variability in shorter texts. In other words, token-level methods benefit from high-frequency watermark signals, whereas SimMark injects at a coarser granularity. As a result, for SimMark, detection performance improves with longer generations, where more sentence-level signal can accumulate.
Figure 4: Detection performance (ROC-AUC ↑) under two adversarial settings using RealNews dataset. Left: Paraphrase+Drop Attack, where random sentences are removed after paraphrasing. SimMark, in nearly all parameter regimes, outperforms other methods under this attack. Right: Paraphrase+Merge Attack, where sentence boundaries (punctuations) are probabilistically replaced with “and” to merge sentences. Although SimMark performs best among sentence-level approaches, UW remains highly robust due to its token-level nature.
Figure 4 presents the ROC-AUC performance of UW (the best token-level method in our experiments), SimMark, $k$-SemStamp, and SemStamp under Paraphrase+Drop and Paraphrase+Merge attacks, evaluated on the RealNews dataset. Under Paraphrase+Drop, across most parameter regimes, SimMark outperforms all other methods, sustaining higher detection performance even as the attack intensity increases. While SimMark shows superior performance among sentence-based methods under Paraphrase+Merge, UW maintains the highest robustness because it operates at the token level.
@@
Regarding sampling efficiency, for BookSum dataset for instance, SimMark required an average of 7.1 samples per sentence from the LLM, compared to $k$-SemStamp and SemStamp, which averaged 13.3 and 20.9 samples, respectively Hou et al. (2024b). This demonstrates that our method not only outperforms these baselines but is also 2-3 times more efficient (see Appendix G for theoretical estimates of this). For a comparison between SimMark and token-level methods in terms of real-world runtime, see Appendix F. Finally, refer to Appendix H for qualitative examples of SimMark.
@@ 5 Conclusion
In this paper, we introduced SimMark, a similarity-based, robust sentence-level watermarking algorithm. Unlike existing approaches, SimMark operates without requiring access to the internals of the model, ensuring compatibility with a wide range of LLMs, including API-only models. By utilizing a pre-trained general-purpose embedding model and integrating a soft counting mechanism, SimMark combines robustness against paraphrasing with applicability to diverse domains. Experimental results show that SimMark outperforms SOTA sentence-level watermarking algorithms in both efficiency and robustness to paraphrasing, representing a step forward in fully semantic watermarking for LLMs.
Limitations
Rejection Sampling Overhead.
The rejection sampling process requires generating multiple candidate sentences until a valid sentence is accepted. Although our method is significantly (2-3 times) more efficient than prior approaches such as SemStamp and $k$-SemStamp, there is still a notable decrease in generation speed due to rejection sampling. Techniques like batch sampling or parallel sampling could potentially mitigate this issue, though at the expense of higher computational resource usage. Future research should focus on optimizing the method to balance efficiency and resource requirements.
Resistance to More Advanced Attacks.
While SimMark demonstrates robustness against paraphrasing attacks, it may not be immune to more sophisticated adversarial transformations. In particular, detection could become less effective when watermarked text is interleaved with or embedded within a larger body of unwatermarked content. Additionally, although reverse engineering the exact watermarking rules is non-trivial, an adversary may attempt a spoofing attack by approximating our setup—for instance, by employing a publicly available embedding model or fitting a PCA model with publicly available datasets. While such attempts may not perfectly replicate the original embedding distribution, they could still pose a threat. We leave a thorough investigation into vulnerabilities and corresponding defences against reverse engineering to future work.
Dependency on Predefined Intervals.
In our experiments, we used consistent, predefined intervals across all datasets and observed consistently strong performance. Notably, we did not observe any noticeable degradation in text quality due to this interval constraint during rejection sampling (as shown in Table 2), likely because the constraint applies only to consecutive sentences. Nonetheless, slight variations in the embeddings similarity distribution of LLM-generated text across different models/datasets may impact watermarking effectiveness. Adaptive strategies for setting these intervals dynamically (or pseudo-randomly) could not only improve performance but also make reverse-engineering the algorithm more difficult.
Ethical Considerations [^16] Generative AI tools, such as ChatGPT, were used to refine this manuscript. The authors retain full responsibility for all content presented in the paper, ensuring adherence to academic integrity and ethical research standards.
Potential Risks.
By enabling robust detection of LLM-generated text, particularly under paraphrasing attacks, SimMark tries to address ethical concerns surrounding the transparency and accountability of AI-generated content. However, like any watermarking algorithm, there are potential risks, such as falsely implicating human authors or adversaries developing more advanced techniques for spoofing attacks or bypassing detection. We acknowledge these limitations and advocate for the responsible deployment of such tools in combination with other verification mechanisms to mitigate these risks and ensure ethical, fair deployment. The primary goal of this work is to advance research in watermarking techniques to support the responsible use of LLMs. We believe that the societal impacts and ethical considerations of our work align with those outlined in Weidinger et al. (2021).
Use of Models and Datasets.
Our research used datasets and pretrained models from the Hugging Face Hub [^17] https://huggingface.co, a public platform hosting machine learning resources under various licenses. We adhered to all license terms and intended usage guidelines for each artifact, which are documented on their individual Hugging Face model or dataset cards cited in the paper. All resources were used solely for research purposes in accordance with their intended use and respective licenses. The datasets employed are publicly available, widely used in prior research, and, to the best of our knowledge, free of personally identifiable information or offensive content. Any outputs generated by our method are intended for academic use, not for real-world or commercial applications, and no personal or sensitive data was processed. Any new artifacts we create (e.g., watermarked samples) are intended solely for academic evaluation, and we do not release any derivative data that violates original licensing terms.
Acknowledgments
@@
Supplemental Materials
Appendix A Aditional Related Work
A.1 Token-Level Watermarking
Zhao et al.’s (2023) UNIGRAM-WATERMARK (UW) builds upon KGW by fixing the red and green lists instead of pseudo-randomly selecting them, proving that, compared to KGW, their method is more robust to paraphrasing and editing Zhao et al. (2023). However, as outlined by Hou et al. (2024a), this algorithm can be reverse-engineered, rendering it impractical for high-stakes, real-world applications (for details about the reverse-engineering procedure, refer to Hou et al. (2024a)).
The Semantic Invariant Robust (SIR) watermark in Liu et al. (2023) is also similar to KGW but is designed to be less sensitive to attacks involving synonym replacement or advanced paraphrasing. SIR achieves this by altering the LLM logits based on the semantics of previously generated tokens, using a semantic embedding model to compute semantic representations, and training a model that adjusts LLM’s logits based on the semantic embeddings of prior tokens Liu et al. (2023).
A.2 Post-Hoc Watermarking
Chang et al.’s (2024) PostMark is a post-hoc watermarking algorithm designed to work without access to model logits, making it compatible with API-only LLMs. It constructs an input-dependent set of candidate words using semantic embeddings and then prompts another LLM (e.g., GPT-4o) to insert these words into the generated text. Detection relies on statistical analysis of the inserted words. While PostMark’s compatibility with black-box LLMs is a strength, the approach is expensive—watermarking 100 tokens is estimated to cost around \$1.2 USD Chang et al. (2024).
Yang et al. (2023) propose another post-hoc method that encodes each word in the text as a binary bit via a Bernoulli distribution ($p=0.5$), embedding the watermark through synonym substitution: words representing bit 0 are replaced with synonyms representing bit 1. Detection is again done via statistical testing. However, this method is fragile: synonyms are not reliably preserved under paraphrasing and often fail to capture subtle contextual meanings, which can noticeably degrade text quality and watermark robustness.
Hao et al. (2025) is a post-hoc watermarking technique similar to Yang et al. (2023) that improves robustness by selecting semantically or syntactically essential words—those less likely to be altered during paraphrasing—as anchor points for embedding. The method uses paraphrase-based lexical substitution to insert watermarks while preserving the original semantics. However, empirical results in Chang et al. (2024) demonstrate that this method is not robust to paraphrasing compared to other methods such as KGW, SemStamp, and PostMark.
Appendix B Experimental Settings
In all combinations of the experiments, following Kirchenbauer et al. (2023), sampling from the LLM was performed with a temperature of 0.7 and a repetition penalty of 1.05, while the minimum and the maximum number of generated tokens were set to 195 and 205, respectively. The maximum number of rejection sampling iterations was set to 100, again to align with the code provided by Hou et al. (2024a, b). However, this setting reflects a trade-off between detection performance and generation speed. Based on our experiments, setting it to 25 achieves strong performance, with higher values offering only marginal improvements (see Appendix E). For token-level watermarking baselines, in cases where results were not directly extracted from Hou et al. (2024a, b), we employed the open-source MarkLLM watermarking framework Pan et al. (2024), with their recommended configurations ($\gamma=0.5,\delta=2$, prefix_length=1, etc.) to run the experiments.
The majority of the experiments, including text generation and detection tasks, were conducted on a workstation equipped with an Intel Core i9 processor, 64GB of RAM, and an Nvidia RTX 3090 GPU with 24GB of VRAM. Some of the experiments involving bigram paraphrasing were performed on compute nodes with an Nvidia V100 GPU with 32GB of VRAM.
Appendix C Averaged Performance Across Datasets
Figure 5: Detection performance averaged across three datasets. Left: ROC-AUC ↑ across different paraphraser variants. Right: TP@FP=5% ↑ under the same settings. These radar plots provide a holistic comparison of all watermarking algorithms across paraphrasing conditions, averaged across datasets. SimMark demonstrates consistently strong robustness, closely matching or outperforming state-of-the-art baselines, particularly under heavier transformations.
Avg. Over Datasets
UW 99.9 / 99.5 / 99.9 98.4 / 82.9 / 95.0 96.9 / 64.3 / 87.3 95.8 / 62.5 / 83.6 95.1 / 56.2 / 80.9 92.3 / 36.9 / 66.2 84.7 / 19.8 / 43.8 93.8 / 53.8 / 76.1
KGW 99.5 / 98.3 / 98.7 95.8 / 86.3 / 91.2 93.4 / 55.5 / 75.3 87.5 / 32.1 / 58.2 86.3 / 26.7 / 53.6 84.8 / 20.6 / 49.9 77.3 / 9.9 / 31.6 87.5 / 38.5 / 60.0
SIR 99.8 / 98.8 / 99.9 92.5 / 69.1 / 82.9 92.9 / 58.5 / 78.4 92.3 / 50.2 / 70.6 93.1 / 51.9 / 72.7 84.1 / 26.5 / 42.6 79.3 / 15.4 / 37.0 89.0 / 45.3 / 64.0
SemStamp 99.5 / 96.6 / 98.0 98.4 / 90.3 / 94.8 97.7 / 85.4 / 91.7 93.9 / 65.1 / 79.2 95.1 / 70.6 / 82.9 86.9 / 42.3 / 57.8 85.3 / 37.5 / 52.8 92.9 / 65.2 / 76.5
$k$-SemStamp 99.8 / 98.6 / 99.1 99.4 / 93.4 / 96.9 99.0 / 90.5 / 95.6 98.1 / 82.5 / 91.7 98.2 / 83.6 / 91.1 93.2 / 60.6 / 77.4 92.3 / 57.4 / 73.8 96.7 / 77.8 / 87.7
Cosine-SimMark (ours) 99.5 / 97.3 / 98.6 99.2 / 92.7 / 97.7 99.1 / 92.7 / 97.5 98.8 / 89.4 / 97.1 98.9 / 88.7 / 97.1 96.9 / 70.2 / 89.4 94.3 / 51.1 / 78.8 97.8 / 80.8 / 92.9
Euclidean-SimMark (ours) 99.9 / 99.1 / 99.5 98.3 / 83.2 / 93.9 98.2 / 80.8 / 92.3 97.1 / 70.5 / 87.7 97.1 / 69.6 / 87.8 97.5 / 76.6 / 90.1 94.8 / 62.5 / 79.5 97.2 / 73.9 / 88.6
Table 3: Detailed average detection performance across the RealNews, BookSum, and Reddit-TIFU datasets under different paraphrasers. Each cell reports ROC-AUC ↑ / TP@1%FP ↑ / TP@5%FP ↑. Higher values indicate better performance across all metrics. Bold and underlined numbers denote the highest and second-highest values, respectively. SimMark consistently ranks among the top-performing methods in robustness across various paraphrasers averaged across datasets.
For a comprehensive comparison across all paraphrasing settings (as shown in Table 1), including ROC-AUC and TP@FP thresholds (1% and 5%), we report the averaged detection performance metrics across datasets in Table 3. To better visualize relative performance trends across paraphrasers, we also present radar plots in Figure 5, showing aggregated TP@FP=5% and ROC-AUC scores (see Figure 3 in the main text for aggregated TP@FP=1% scores). These visualizations complement the table by providing an intuitive view of robustness across settings. Together, these summaries highlight that SimMark consistently outperforms or matches the baseline methods across diverse conditions.
Appendix D Additional Experimental Results
To demonstrate the model-agnostic nature of SimMark, we applied our algorithm to the Gemma3-4B model [^1] We employed google/gemma-3-4b (4B) model. Team et al. (2025). We evaluated both Cosine-SimMark and Euclidean-SimMark under different paraphrasing models across the same three datasets as before: RealNews subset of C4, BookSum, and Reddit-TIFU. The effectiveness and robustness of watermarking techniques can depend heavily on the characteristics of the underlying LLM and the nature of the generated text. To maintain consistency and reliability across experiments, we made the following modifications:
  • Predefined Interval Adjustment: The sentences’ embedding similarity distribution under Gemma3-4B differed from those in OPT-1.3B, requiring new intervals. We set the predefined interval to [0.86, 0.90] for cosine similarity (without PCA), and [0.11, 0.16] for Euclidean distance (with PCA).
  • Threshold Transferability: In contrast to our earlier experiments—where the detection threshold $\beta$ was determined per dataset—we fixed $\beta$ across all datasets in these experiments. Specifically, we determined the threshold using only non-watermarked data from the BookSum dataset, and then applied it across all three datasets without modification. This approach simulates a more realistic setting where the detector is calibrated on a single corpus but expected to generalize to others. The results demonstrate that our method maintains high detection performance even under this general configuration.
  • Longer Generations: Since Gemma3-4B tends to generate longer sentences compared to OPT-1.3B model that we employed earlier, we increased the number of generated tokens from 200 to 300 to ensure a sufficient number of sentences for reliable hypothesis testing.
Table 4 reports the detection performance in terms of ROC-AUC ↑ / TP@1%FP ↑ / TP@5%FP ↑, for each setting (↑: higher is better). Across all datasets and paraphrasing scenarios, SimMark remains highly effective, with both cosine similarity and Euclidean distance variants maintaining strong ROC-AUC and TP rates. These results affirm that SimMark maintains its performance across different LLM families (i.e., OPT and Gemma3) and datasets/domains, further validating the general applicability of our proposed algorithm.
Dataset Method No Paraphrase Pegasus Pegasus-Bigram Parrot Parrot-Bigram Avg. Paraphrased
RN
Cosine-SimMark 99.5 / 96.2 / 97.9 93.5 / 57.9 / 75.6 93.0 / 54.9 / 73.9 92.3 / 50.1 / 71.7 92.1 / 49.2 / 72.8 92.7 / 53.0 / 73.5
Euclidean-SimMark [^7] PCA is applied. 99.5 / 97.4 / 98.2 93.2 / 65.3 / 79.5 91.8 / 61.8 / 78.5 91.9 / 58.0 / 73.9 91.5 / 58.7 / 72.8 92.1 / 61.0 / 76.2
BS
Cosine-SimMark 99.9 / 99.6 / 99.8 97.7 / 75.8 / 90.5 97.4 / 73.3 / 90.0 97.0 / 67.8 / 87.6 96.9 / 67.7 / 86.7 97.2 / 71.1 / 88.7
Euclidean-SimMark [^7] PCA is applied. 99.9 / 99.7 / 99.9 97.6 / 82.0 / 91.4 97.4 / 77.9 / 89.9 97.4 / 78.0 / 91.2 91.5 / 58.7 / 72.8 96.0 / 74.2 / 86.3
TIFU
Cosine-SimMark 99.8 / 99.1 / 99.5 97.4 / 78.8 / 91.5 97.0 / 76.6 / 89.3 89.4 / 32.8 / 61.7 90.7 / 36.6 / 63.8 93.6 / 56.2 / 76.6
Euclidean-SimMark [^7] PCA is applied. 99.6 / 98.8 / 99.1 97.3 / 82.0 / 91.4 96.9 / 81.0 / 90.1 92.2 / 53.6 / 73.9 92.8 / 58.4 / 76.0 94.8 / 68.8 / 82.8
Table 4: Performance of SimMark using Gemma3-4B model across paraphrasers and datasets (RealNews denoted as RN, BookSum denoted as BS, and Reddit-TIFU denoted as TIFU). Each cell reports AUC $\uparrow$ / TP@FP=1% $\uparrow$ / TP@FP=5% $\uparrow$ (↑: higher is better). The results demonstrate that SimMark maintains strong performance across different datasets and paraphrasing conditions, highlighting its robustness and model-agnostic nature.
Appendix E Effect of Sampling Budget on Detection Performance
To better understand the trade-off between generation speed and detection performance, we analyze the impact of the max_trials hyperparameter, which defines the upper limit on the number of rejection sampling iterations during watermark injection. While we set this value to 100 in our main experiments (to align with prior works of Hou et al. (2024a, b)), it is important to examine whether such a large value is necessary.
Figure 6 shows two evaluation metrics—ROC-AUC ↑ and TP@FP=1% ↑ (↑: higher is better)—on the RealNews dataset for cosine-SimMark and OPT-1.3B model under different values of max_trials. As shown in the plots, performance improves significantly when increasing max_trials from 5 to 25, but plateaus thereafter. In particular, both ROC-AUC and TP@FP=1% show diminishing returns beyond 25 trials, indicating that additional sampling brings little performance gain.
These results suggest that setting max_trials to 25 achieves a good balance between robustness and efficiency, and using larger values (e.g., 100) is not strictly necessary in practice. These findings, together with the average sampling statistics reported in the main paper (e.g., 7.1 samples per sentence), highlight SimMark’s ability to balance robustness with generation speed compared to SemStamp (20.9 samples per sentence) and $k$-SemStamp (13.3 samples per sentence).
Figure 6: Impact of the maximum number of rejection sampling trials on detection performance. Increasing max_trials improves both ROC-AUC ↑ and TP@1%FP ↑ (↑: higher is better), but the improvement plateaus around 25. Results are reported on the RealNews dataset using cosine-SimMark and OPT-1.3B model.
Appendix F Empirical Analysis of Generation Efficiency
We experimentally evaluate the generation-time latency introduced by SimMark. Using 100 samples from the BookSum dataset, we measure the per-sentence generation time under various configurations of max_trials, using OPT-1.3B model. On our hardware (see Appendix B), with max_trials = 100 (i.e., maximum number of rejection sampling), cosine-SimMark yields an average latency of 1.33 seconds per sentence—corresponding to a 7.1× slowdown compared to the unwatermarked baseline average (0.188 sec/sentence). This matches our reported average sampling rate of 7.1 candidates per sentence (see Section 4).
For reference, token-level methods such as UW and SIR yield 0.182 and 0.377 seconds per sentence, respectively (i.e., around 1× and 2× slowdown). This highlights the key trade-off between efficiency and robustness: while token-level methods are faster, SimMark offers less impact on text quality and stronger resistance to paraphrasing.
Method Latency Slowdown
Baseline (no watermark) 0.188 sec/sent 1.0×
SimMark (max_trials=25) 0.977 sec/sent 5.2×
SimMark (max_trials=50) 1.114 sec/sent 5.9×
SimMark (max_trials=100) 1.330 sec/sent 7.1×
UW (token-level) 0.182 sec/sent $\sim$1.0×
SIR (token-level) 0.377 sec/sent 2.0×
Table 5: Latency comparison of different watermarking methods (in seconds per sentence) and their slowdowns relative to unwatermarked generation. SimMark’s higher overhead stems from rejection sampling but remains practical—especially at lower max_trials—and can be further reduced with efficient decoding backends (e.g., vLLM).
In table 5, we also evaluate SimMark under smaller max_trials values for practical trade-offs, as performance remains strong even with a smaller sampling budget (see Appendix E).
While SimMark introduces overhead due to rejection sampling, newer decoding backends (e.g., vLLM Kwon et al. (2023)) can reduce this cost substantially via prompt-prefix reuse and KV cache optimization. However, our current benchmarks were conducted using the Hugging Face Transformers library Wolf et al. (2020) without these optimizations.
Although we were unable to directly benchmark SemStamp and $k$-SemStamp due to technical issues, their substantially higher sampling requirements suggest that they would incur 2–3× greater overhead than SimMark. It is worth noting that a generation latency of about 1 second per sentence remains below the average human reading speed (200–250 wpm, or roughly 3–6 seconds per sentence). In summary, we believe SimMark offers a reasonable trade-off between detection robustness and runtime overhead, maintaining practicality for many real-world applications even in long-form scenarios.
Appendix G Theoretical Analysis of Sampling Efficiency
The average number of samples required to generate a valid sentence is influenced by the chosen interval $[a,b]$. To provide further insights into this relationship, we estimated the area under the curve (AUC) of the embedding similarity distribution for an unwatermarked LLM (OPT-1.3B). For instance, for the interval $[0.68,0.76]$ (for cosine-SimMark), the estimated AUC is approximately $0.194$.
Using the mean of the geometric distribution, which is given by $\frac{1}{p}$, where $p$ is the probability of success (in this case, the probability of falling within the interval), this translates to an expected average of $\frac{1}{0.194}\approx 5.1$ samples per valid sentence. This estimate is on par with the experimental results reported in the main paper. The AUC estimates were computed using the binning technique, as described in Appendix L.
This analysis underscores the importance of carefully selecting the interval $[a,b]$, as narrower intervals may increase the number of required samples, leading to reduced sampling efficiency but better performance, while broader intervals may compromise the effectiveness of the watermark. By understanding the interplay between the interval choice and sampling efficiency, we can better optimize SimMark’s performance.
Appendix H Examples of Watermarked Text